### v2

# DualToken: Towards Unifying Visual Understanding and Generation with Dual Visual Vocabularies

Abstract
The differing representation spaces required for visual understanding and generation pose a challenge in unifying them within the autoregressive paradigm of large language models. A vision tokenizer trained for reconstruction excels at capturing low-level perceptual details, making it well-suited for visual generation but lacking high-level semantic representations for understanding tasks. Conversely, a vision encoder trained via contrastive learning aligns well with language but struggles to decode back into the pixel space for generation tasks. To bridge this gap, we propose DualToken, a method that unifies representations for both understanding and generation within a single tokenizer. However, directly integrating reconstruction and semantic objectives in a single tokenizer creates conflicts, leading to degraded performance in both reconstruction quality and semantic performance. Instead of forcing a single codebook to handle both semantic and perceptual information, DualToken disentangles them by introducing separate codebooks for high and low-level features, effectively transforming their inherent conflict into a synergistic relationship. As a result, DualToken achieves state-of-the-art performance in both reconstruction and semantic tasks while demonstrating remarkable effectiveness in downstream MLLM understanding and generation tasks. Notably, we also show that DualToken, as a unified tokenizer, surpasses the naive combination of two distinct types vision encoders, providing superior performance within a unified MLLM. The code and models will be available at https://github.com/songweii/DualToken.
1 Introduction
Figure 1: Comparison with state-of-the-art vision encoders. (Left) We compare zero-shot classification accuracy and reconstruction FID on ImageNet-1K(val) across baseline methods and DualToken. DualToken achieves results comparable to or surpassing both semantic-only and reconstruction-only methods in both tasks. (Right) Reconstruction results of VILA-U and DualToken, our DualToken significantly outperforms VILA-U, which suffers from severe distortion and blurriness.
Unifying visual understanding and generation within the autoregressive paradigm of large language models (LLMs) has become a current research hotspot, giving rise to representative works like CM3leon [54], Chameleon [40], Emu3 [45] and VILA-U [47]. To achieve multimodal autoregressive generation, these unified models require a visual tokenizer to discretize visual inputs and a corresponding detokenizer to map tokens back to the pixel space.
Early methods [54, 40, 45] directly used the encoder and decoder of VQVAE as the tokenizer and detokenizer. However, although these methods successfully demonstrated the feasibility of unifying visual understanding and generation within the autoregressive paradigm, their understanding capabilities are typically lacking compared to traditional multimodal large language models (MLLMs) that specialize in understanding tasks [29, 12, 55, 58, 38, 20, 31, 32]. We argue that this decline in visual understanding performance stems from the insufficient visual representations—traditional VQVAE pre-training focuses on reconstruction tasks, making its embedding space rich in low-level visual information, but lack higher-level semantic knowledge. In contrast, MLLMs designed for understanding tasks [26, 27, 28, 2, 44, 30, 48, 23, 24] typically utilize CLIP-family encoders [35, 56, 1] to extract visual features. Since these encoders are pretrained with language alignment, their representations inherently capture high-level semantic information, making them more suitable for downstream visual understanding tasks in MLLMs.
To fully leverage the text-aligned semantic representations of CLIP, a natural approach is to quantize the features of a CLIP encoder and train a corresponding decoder for image reconstruction [47]. Specifically, this requires the visual tokenizer to learn reconstruction for downstream generation tasks while preserving its semantic capabilities as much as possible [47]. However, as shown in Fig.1 and Table1, directly combining reconstruction and semantic objectives often leads to severe distortions and blurriness in reconstruction tasks, along with a noticeable decline in semantic metrics such as zero-shot classification and image-text retrieval, compared to its original pretrained model [56]. This degradation, as discussed in [47], reflects the inherently conflict between the two training objectives, ultimately limiting both the quality of downstream image generation tasks and the performance of multimodal understanding tasks.
Table 1: Comparison to state-of-the-art visual encoders or tokenizers. For semantic metrics, we measured ImageNet zero-shot classification accuracy, as well as Text-to-Image/Image-to-Text retrieval (R@1) on Flickr8K. For reconstruction, we measured reconstruction FID (rFID), PSNR, and SSIM on ImageNet-1K (Val). Our method not only outperforms VILA-U and achieves performance comparable to the state-of-the-art SigLIP ViT-SO400M model in semantic metrics, but also mitigates the structural distortion and blurriness issues faced by VILA-U during reconstruction. Our method also surpasses dedicated models such as MoVQGAN in reconstruction metrics, reaching state-of-the-art performance.
Vision encoder (tokenizer) Semantic Reconstruction
Zero-Shot T2I(R@1) I2T(R@1) rFID  PSNR SSIM
VQGAN [11] 4.98 - -
MoVQGAN [59] 1.12 22.42 0.673
RQ-VAE [19] 2.69 - -
ViT-VQGAN [52] 1.55 - -
Open-MAGVIT2 [53, 33] 1.17 21.90 -
SBER-MoVQGAN [37] 0.68 27.04 0.741
FQGAN-Dual ($\downarrow$16) [3] 0.94 22.02 -
FQGAN-Triple ($\downarrow$16) [3] 0.76 22.73 -
clip-ViT-L-14-336 [35] 76.6 21.15 21.50
siglip-so400m-14-384 [56] 82.1 21.70 21.60
VILA-U [47] 73.3 10.02 11.15 1.80 3.43 0.489
VILA-U [47] (SO400M) 78.0 - - 1.25 - -
DualToken ($\downarrow$16) (ours) 81.6 21.05 21.55 0.54 23.56 0.742
To decouple the two conflicting objectives effectively, we propose using two sets of visual vocabularies—High-Level (Semantic) and Low-Level (Perceptual)—to tokenize semantic and texture features separately. Specifically, similar to the hierarchical structure of the human vision system [16], the shallow-layer features of a ViT focus on perceptual-level texture information, while high-level semantic representations emerge in the deeper layers [6]. We argue that this inherent property of ViT should be fully leveraged: using shallow-layer features for reconstruction and deep-layer features for semantic learning, thereby obtaining the texture and semantic codebook simultaneously.
Surprisingly, this hierarchical decoupling not only resolves the conflict between the two objectives but also enables the semantic learning objective to enhance low-level reconstruction. Moreover, training the shallow-layer reconstruction task does not compromise the model’s original semantic capabilities, even without additional contrastive learning stages [35, 47]. Building upon this, we further demonstrate how a multimodal large language model (MLLM) can effectively utilize the dual visual vocabularies to achieve unified vision understanding and generation.
2 Related Work
Unified Multimodal Large Language Models A common strategy for integrating visual understanding and generation within a single MLLM is to externally connect an LLM with a Diffusion Model [39, 9, 13, 14]. However, pure autoregressive (AR) architectures offer a more elegant, fully end-to-end solution by unifying both tasks within the same autoregressive framework. Representative models like Chameleon [54, 40] and Emu3 [45], have demonstrated the feasibility of jointly modeling vision and language through a unified next-token prediction objective. Specifically, visual inputs are first discretized into visual tokens with a vision tokenizer. These visual tokens are then interleaved with text tokens to construct a multimodal sequence. However, pure AR architectures like Chameleon introduce generative capabilities at the cost of significantly weaker visual understanding compared to models specifically designed for understanding tasks. An empirical explanation for this [47, 50] is that models like Chameleon typically utilize VQVAE [43, 11] encoders as their vision tokenizers. Since VQVAE is trained with a reconstruction objective, its encoder primarily extracts low-level visual features optimized for generation rather than the high-level semantic representations required for vision-language understanding. Therefore, improving the vision-language understanding performance of AR-based models necessitates a vision tokenizer that effectively captures both low-level representations for generation and high-level semantic representations for understanding.
Recent research has actively explored solutions in this direction. VILA-U [47] and MUSE-VL [50] strive to build a unified tokenizer by jointly training on both reconstruction and semantic objectives. However, due to the inherent disparity between semantic and texture features, they struggle to strike an optimal balance between the two objectives, resulting in subpar performance in both tasks. As discussed in FQGAN [3], decomposing the codebook in a divide-and-conquer manner may offer a more fundamental solution to this conflict. TokenFlow [34] employs separate codebooks with a shared-mapping mechanism. However, key differences set our approach apart. First, TokenFlow relies on distinct vision towers to extract semantic and low-level features, rather than leveraging a unified architecture. Second, the shared IDs obtained through the shared-mapping mechanism may not be the optimal matches for either semantics or texture, potentially introducing additional losses in both domains. TokenFlow has yet to demonstrate its effectiveness in a unified MLLM setting that supports both generation and understanding; instead, it trains separate models dedicated to each task.
3 Method
Figure 2: Overview of our unified vision tokenizer. Given input images the features extracted by the vision encoder are discretized using residual quantization. Then the discrete vision features are meanwhile put into the vision decoder to reconstruct images and used to perform the text-image alignment. During this process, the reconstruction loss and contrastive loss are computed to update the vision tower, endowing it to produce discrete visual features with text alignment.
This section formally introduces the design of our unified tokenizer and explains how its dual visual codebooks are utilized within the next-token prediction paradigm of LLMs for unified multimodal understanding and generation.
3.1 Motivation
As discussed in [34], CLIP-based encoders cluster images based on semantic similarity, whereas VQVAE-based encoders group images by low-level attributes such as color and texture. This suggests that encoders trained for reconstruction primarily extract low-level perceptual information, while CLIP-family encoders, pretrained with language alignment, inherently capture high-level semantic information. We argue that this difference in representation space plays a crucial role in downstream MLLM performance.
To validate this viewpoint, we started by a preliminary experiment following the LLaVA-1.5 pipeline [27]. As shown in Table 2, compared to the original SigLIP model, encoders trained with reconstruction objective exhibit a significant drop in downstream MLLM vision-language understanding performance, validating that high-level semantic features are more critical for visual reasoning in MLLMs than low-level perceptual features. However, to achieve both visual understanding and generation within a single MLLM, it is essential to decode the visual tokens back into pixel space as accurately as possible. However, since the SigLIP encoder focuses on high-level semantic information rather than texture details, simply discretizing its features and training a decoder without tuning the encoder results in poor image reconstruction quality. Therefore, proposing a unified tokenizer is crucial to enable high-quality visual understanding and generation within a singe MLLM.
Table 2: Downstream visual understanding performance with different vision encoders within the LLaVA-1.5 framework. ViT-SO400M (pretrained) refers to the original pretrained siglip-so400m-14-384 model [1], while ViT-SO400M (recon.) refers to the encoder that follows the same architecture but is trained solely for reconstruction from scratch, controlling for factors like model size and architecture.
Vision Encoder MMB MME SEED VQAv2 Zero-Shot rFID
ViT-SO400M (pretrained) 61.8 1492.9 56.8 78.5 82.1
ViT-SO400M (recon.) 36.2 822.4 30.6 47.5 0.96
SBER-MoVQGAN-270M 35.8 792.0 34.1 45.2 0.68
3.2 Unified Vision Tokenizer with Dual Codebooks
To build a unified tokenizer, we started with the simplest approach, where we directly combine the reconstruction loss and semantic loss to optimize the entire vision tower and use a single visual vocabulary to tokenize its feature, similar to VILA-U [47]. Specifically, as illustrated in Fig.2 (left), we initialize the vision encoder with pretrained weights from siglip-so400m-patch14-384 [1] to ensure strong text-image alignment. Then the semantic loss is computed between the deeper-layer features of the model and its initial state to constrain the model from losing its semantic capability.
However, as shown in Table.3(a), this straightforward approach leads to a clear conflict between the two objectives. On one hand, although the semantic loss is applied to preserve the model’s original semantic representation capabilities, achieving this objective proves difficult, as semantic performance metrics show a significant decline compared to the original model, reflecting the disruption caused by the reconstruction training objective on semantic capabilities. On the other hand, the model also struggles to achieve satisfactory reconstruction quality, often resulting in distortions and blurriness in the reconstructed images (Fig.2).
To address this, we decouple the learning of the reconstruction and semantic objectives through a hierarchical approach, as shown in Fig.2 (right). Specifically, reconstruction loss is applied to supervise the shallow layers (1-6) of the vision tower, while semantic loss is applied to the deeper layers (layer 26 and the pooling head). Features from the shallow and deep layers are then discretized separately via residual vector quantization [19], resulting in low-level and high-level visual vocabularies.
We utilize an Exponential Moving Average (EMA) strategy [36, 42] to update the codebook. To enhance codebook utilization, we implement a restart strategy where a cluster within the codebook is randomly replaced with an input from the current batch if it remains unused for a certain number of steps. To ensure the encoder outputs align closely with the codebook entries, we utilize a Vector Quantization (VQ) commitment loss, which is defined as:
$\mathcal{L}_{c}=\|z-\text{quantize}(z)\|_{2}^{2}$ (1)
Consequently, the total loss is formulated as a weighted sum of reconstruction loss, semantic loss, and VQ commitment loss
$\mathcal{L}_{total}=\lambda_{1}\cdot\mathcal{L}_{recon}+\lambda_{2}\cdot\mathcal{L}_{sem}+\lambda_{3}\cdot(\mathcal{L}_{c1}+\mathcal{L}_{c1})$ (2)
where the reconstruction loss is the combination of pixel-wise $L2$ loss [10], LPIPS loss [57] and adversarial loss [18] for reconstructing an input image:
$\mathcal{L}_{recon}=\|\hat{x}-x\|_{2}^{2}+\lambda_{p}\mathcal{L}_{\text{LPIPS}}(\hat{x},x)+\lambda_{g}\mathcal{L}_{\text{G}}(\hat{x})$ (3)
while the semantic loss is simply computed as the $L_{2}$ distance between the model’s 26th-layer feature representation $F$ and its corresponding initial state $F_{0}$
$\mathcal{L}_{sem}=\|F-F_{0}\|_{2}^{2}$ (4)
Table 3: DualToken transforms the conflict between reconstruction and semantic objectives into a synergistic relationship. Directly combining the two objectives leads to a drastic decline in reconstruction performance (a vs. b). However, incorporating reconstruction and semantic losses hierarchically results in better reconstruction performance compared to using reconstruction alone as the target (d vs. c). We highlight our method in the last row.
# Exp. Loss Type (layer) Zero-Shot Acc. Reconstruction
rFID  PSNR SSIM
Initial State: 82.1
    (a) Recon. (26) + Sem. (26) 75.2 4.36 11.85 0.579
    (b) Recon. (26) 0.71 21.88 0.640
    (c) Recon. (6) 0.73 22.12 0.680
    (d) Recon. (6) + Sem. (26) 81.6 0.54 23.56 0.742
Interestingly, as shown in Table.3 (d), even without adding an additional contrastive learning phase to enhance semantic capabilities and relying solely on a simple $L2$ loss to constrain the semantic representation, incorporating a reconstruction learning objective in our hierarchical learning strategy causes minimal damage to the model’s semantic ability. More intriguingly, as shown in Table.3 (b), (c), and (d), compared to training solely for reconstruction, learning the semantic objective in the deeper layers actually enhances the reconstruction task in the shallow layers, successfully transforming the conflict between semantic and reconstruction objectives into a positive relationship.
3.3 Unifying Understanding and Generation
Figure 3: An overview of our framework utilizing dual visual codebooks for unified visual understanding and generation. (a) Directly using VQGAN and SigLIP to separately acquire high-level (semantic) and low-level (pixel) visual codebooks. (b) Our approach: decoupling high-level and low-level visual codebooks within a unified vision tokenizer. The image is converted into low-level visual tokens (green) and text-aligned semantic tokens (red). (c) To model both textual and visual content within the autoregressive paradigm of LLMs, the pixel and semantic visual tokens are first concatenated along their embedding dimension to form unified visual tokens (yellow). These unified visual tokens are then concatenated with text tokens to construct a multimodal token sequence. The model is trained to predict the next token for both visual and textual tokens. Specifically, the high-level and low-level visual tokens are processed by independent visual heads (Pixel head and Semantic head), each comprising a depth transformer (3 layers with a depth of 8) and 8 classification heads to predict the residuals of the corresponding visual token at different depths. During inference, the generated low-level tokens are decoded by our visual decoder to reconstruct the visual content.
In this section, we demonstrate how to integrate the dual visual codebooks of DualToken within a unified MLLM. As illustrated in Fig. 3 (c), to model both textual and visual content under the autoregressive paradigm of LLMs, the pixel and semantic visual tokens are first concatenated along their embedding dimension to form unified visual tokens. These unified visual tokens are then concatenated with text tokens to construct a multimodal token sequence. Then the model is trained in an autoregressive manner to predict the next token for both visual and textual tokens. For simplicity, we define the language vocabulary of our MLLM as a finite set $\mathcal{X}=\{x_{1},x_{2},...,x_{n_{1}}\}$, while the low-level and high-level visual vocabulary as $\mathcal{Y}=\{y_{1},y_{2},...,y_{n_{2}}\}$ and $\mathcal{Z}=\{z_{1},z_{2},...,z_{n_{3}}\}$, where $n_{1}$, $n_{2}$, and $n_{3}$ represent the vocabulary sizes for language tokens, low-level visual tokens, and high-level visual tokens, respectively.
For visual tokens, since residual quantization introduces a depth-stacked structure of codes at each visual position $p$, we implement our visual heads based on the depth transformer from RQ-VAE [19]. Unlike the original depth transformer, which employs a single head to predict logits across all depths, we introduce separate classification heads to compute the logits for residuals at each corresponding depth [21]. Specifically, the high-level semantic tokens and low-level pixel tokens are processed by independent visual heads—the pixel head and the semantic head—as shown in Fig.3. Both heads share the same structure, comprising three layers of depth transformers (each with a depth of 8) and eight classification heads.
Given the LLM hidden state $h_{p}$ for visual tokens at position $p$, our depth transformer autoregressively predicts D residual tokens ($r_{p1}$, $r_{p2}$, …, $r_{pD}$). For $d>1$, the input to the depth transformer at depth d, denoted as $I_{pd}$, is defined as the sum of the token embeddings of up to depth $d-1$
$I_{pd}=\sum_{d^{\prime}=1}^{d-1}\mathbf{e}(r_{pd^{\prime}}),$ (5)
where $r\in\mathcal{Y}$ for the pixel head and $r\in\mathcal{Z}$ for the semantic head. The initial input at depth 1 is given by $I_{p1}=h_{p}$. This formulation ensures that the depth transformer incrementally refines the predicted feature representation by leveraging previous estimations up to depth $d-1$. Consequently, the overall negative log-likelihood loss for the entire multimodal sequence of length $N$ is defined as
$\mathcal{L}_{\text{NTP}}=-\sum_{i=1}^{N}\mathcal{P}_{i}$ (6)
where,
$\mathcal{P}_{i}=\log P\left(x_{i}|x_{<i}\right)$ (7)
if a text token appears at position $i$, and
$\mathcal{P}_{i}=\sum_{d=1}^{D}~{}[\log P\left(y_{id}|y_{i,<d}\right)+\log P\left(z_{id}|z_{i,<d}\right)]$ (8)
if visual tokens appears at position $i$. During multimodal pretraining, the weights of the depth transformers are randomly initialized and trained jointly with the LLM. During inference, only the low-level tokens are utilized by our visual decoder to reconstruct the visual content.
4 Experiments
In this section, we present a comprehensive set of experiments to assess our method across a range of visual understanding and generation tasks. We begin by detailing our experimental setup. Next, we analyze the performance of our unified vision tokenizer. Finally, we benchmark our approach against leading MLLMs, showcasing its strengths in both visual understanding and generation. However, please note that this research is still ongoing, and the final results, complete metrics, and additional technical details may be updated before the final release.
Figure 4: Visual generation results with DualToken. (Left) Our DualToken can generate high-quality images given text input. (Right) Following the pipeline introduced in Fig. 3 (a), we experimented with using the codebook of SBER-MoVQGAN as the low-level vocabulary and the codebook of a VQ-processed SigLIP as the high-level vocabulary, while keeping the same method and training data for downstream MLLM training. However, this straightforward approach results in significantly poorer image generation performance.
4.1 Experimental Setup
We utilize Qwen-2.5-3B [51] as the base language model and adopt the pretrained weights from the SigLIP-SO400M-patch14-384 [1] for our visual tokenizer. All images are resized to $384\times 384$ and transformed into $27\times 27\times 8$ semantic or pixel tokens, with a residual depth of $D=8$. Our vision tokenizer is trained on CC12M [4] and evaluated for zero-shot classification and reconstruction performance on ImageNet [8]. We evaluate our model against widely used vision-language understanding benchmarks, including VQAv2 [15], POPE [22], MME [12], SEED-IMG [20], MMBench [29], and MM-Vet [55]. For visual generation, we apply classifier-free guidance [17] with a CFG value of 3 to enhance the quality of generated outputs.
4.2 Vision Tokenizer
For evaluating the semantic capabilities of our unified vision tokenizer, we report the Top-1 accuracy for zero-shot image classification on ImageNet-1K (validation set), along with text-to-image and image-to-text retrieval performance (R@1) on Flickr8K. As shown in Table 1, our DualToken significantly outperforms VILA-U in both classification and retrieval tasks, while also surpassing the dedicated CLIP-L-14-336 model in zero-shot image classification. Notably, thanks to our hierarchical decoupling approach, DualToken achieves semantic performance on par with the state-of-the-art SigLIP ViT-SO400M-14-384 model, without requiring any additional stages specifically designed to enhance semantic capabilities. We believe that incorporating an additional contrastive learning stage—where the shallow layers responsible for reconstruction are frozen while only the deeper layers are optimized for semantic objective—could further enhance the model’s semantic performance.
To evaluate reconstruction capability, we measured reconstruction FID (rFID), PSNR, and SSIM on the ImageNet-1K validation set. Our DualToken achieves the highest structural similarity and the lowest rFID among various state-of-the-art dedicated methods, including Open-MAGVIT2 [33] and SBER-MoVQGAN [37]. This demonstrates that our method effectively mitigates the structural distortion and blurriness issues encountered by VILA-U during reconstruction.
Table 4: Controlled experiment on various vision-language understanding benchmarks. We evaluate different vision encoders/tokenizers, including siglip-so400m-14-384, VILA-U, and DualToken within the LLaVA-1.5 framework. MMB denotes MMBench-dev.
Vision Encoder MMB VQAv2 POPE MME SEED MMVet
siglip-so400-patch14-384 61.8 78.5 85.9 1492.9 56.8 36.0
VILA-U 56.3(-5.5) 74.1(-4.4) 83.5(-2.4) 1301.4(-191.5) 52.2(-4.6) 25.9(-10.1)
DualToken (sem.+pcpt.) 64.5(+2.7) 78.3(-0.2) 86.1(+0.2) 1589.6(+96.7) 58.9(+2.1) 37.8(+1.8)
DualToken (sem.) 61.2(-0.6) 77.8(-0.7) 84.8(-1.1) 1495.8(+2.9) 57.2(+0.4) 35.4(-0.6)
Table 5: Evaluation on multimodal understanding benchmarks. Our DualToken (3B) demonstrates strong performance compared to other unified models and achieves results comparable to dedicated understanding models like LLaVA-NeXT and ShareGPT4V. Note that the latest version is still in training, and the metrics may be updated before the final release.
Type Method # LLM Params POPE MMBench SEED MMMU MMVet MathVista MME
InstructBLIP [7] 7B - 36.0 58.8 30.6 26.2 24.4 1137.1
LLaVA-Phi [60] 2.7B 85.0 59.8 - - 28.9 - 1335.1
LLaVA-1.5 [27] 7B 85.9 64.3 58.6 35.4 31.1 27.4 1510.7
Und. LLaVA-NeXT [28] 7B 86.5 67.4 70.2 35.8 43.9 34.6 1519.0
LLaVA-NeXT [28] 34B 87.7 79.3 75.9 51.1 57.4 46.5 1631.0
ShareGPT4V [5] 7B - 68.8 69.7 37.2 37.6 26.5 1567.4
VILA [25] 7B 85.5 68.9 61.1 - 34.9 - 1533.0
DreamLLM [9] 7B - 58.2 - - 36.6 - -
SEEDLLaMA [13] 7B - 45.8 51.5 - - - -
Chameleon [40] 7B - 31.1 - 22.4 8.3 - -
Emu3 [45] 8B 85.2 58.5 68.2 31.6 - - -
Uni. Show-o [49] 1.5B 73.8 - - 25.1 - - 948.4
Janus [46] 1.5B 87.0 69.4 63.7 30.5 34.3 - 1338.0
MUSE-VL [50] 7B - 72.1 70.0 42.3 - 52.5 1480.9
VILA-U [47] 7B 85.8 - 59.0 - 33.5 - 1401.8
DualToken (ours) 3B 86.0 70.9 70.2 38.6 32.5 46.5 1489.2
4.3 Downstream Performance
Visual Understanding Model.
Before formally presenting the performance of our unified model, we first conducted a controlled experiment to validate the effectiveness of our vision tokenizer in downstream MLLM visual understanding tasks. To ensure a fair comparison by controlling factors such as training data, model size, and architecture, we evaluate the downstream visual understanding performance of DualToken within the LLaVA-1.5 [27] framework. Specifically, we replace the vision encoder of LLaVA-1.5 with DualToken, while strictly adhering to its training data and using LLaMA-2-7B [41] as the foundational LLM. As shown in Table 4, our DualToken, as a discrete unified vision tokenizer, outperforms VILA-U and even surpasses the original continuous SigLIP model when used as the vision encoder. More interestingly, we conducted separate experiments using only the semantic tokens (sem.), extracted from the 26th layer, as well as a combination of semantic tokens and pixel tokens (sem.+pcpt), concatenated along the embedding dimension as visual input. Compared to using semantic tokens alone, jointly leveraging semantic and pixel tokens generally leads to better performance across various visual reasoning benchmarks, such as MMBench [29] and MME [12]. This highlights that low-level texture features are not only essential for generation tasks but also contribute positively to enhancing the model’s visual understanding capabilities.
Unified Model for Generation and Understanding.
We further implemented our unified MLLM framework for both visual understanding and generation based on the method introduced in Sec.3.3 and Fig.3, where separate visual heads predict semantic and pixel tokens. As shown in Table 5, our DualToken (3B) demonstrates strong performance compared to other unified models and achieves results comparable to dedicated understanding models like LLaVA-NeXT and ShareGPT4V. Meanwhile, as illustrated in Fig. 4, our unified model can generate visually compelling content from text input. The generated images exhibit remarkable alignment with the text, even for long and complex prompts. Thanks to the high reconstruction quality of DualToken, the generated images are rich in detail and structurally realistic, accurately capturing fine textures such as animal fur, water waves, and other intricate patterns.
To further answer a basic question: why do we need to obtain dual visual vocabularies within a unified tokenizer rather than simply combining existing specialized encoders? We experimented with using the codebook of SBER-MoVQGAN as the low-level vocabulary and the codebook of a VQ-processed SigLIP as the high-level vocabulary, while keeping the same method and training data for downstream MLLM training. As shown in Fig. 4 (right), this straightforward approach results in significantly poorer image generation performance, further demonstrating the importance of obtaining dual visual vocabularies within a unified visual tokenizer. It is also worth noting that this simple combination approach differs from Janus [46], whose semantic encoder operates continuously without tokenization and only serves as the encoder for understanding tasks, without being involved in the visual generation process.
5 Conclusion
In summary, our contributions are threefold: (i) Decoupling Reconstruction and Semantic Objectives: We successfully disentangle reconstruction and semantic learning objectives through a hierarchical approach, transforming their inherent conflict into a beneficial relationship. (ii) Dual Visual Codebooks for Enhanced Understanding and Generation: We demonstrate that employing dual visual codebooks outperforms single-codebook approaches in both understanding and generation tasks. (iii) A Unified Model for Vision Understanding and Generation: We propose a viable paradigm for unifying vision understanding and generation using dual visual codebooks. However, this work represents only a baseline implementation of this paradigm, leaving ample room for further exploration. Notably, prior unified models for understanding and generation have not demonstrated clear mutual reinforcement between these tasks. A key direction is to investigate the potential of fully leveraging dual visual codebooks to achieve genuine synergy between visual understanding and generation.
Important Note
Please note that this research is still in progress. The final results, complete metrics, and additional technical details are scheduled to be updated in due course.